OpenStudy (anonymous):

What is the factorization of the trinomial below? 3x2 + 27x + 60

OpenStudy (unklerhaukus):

\[3x^2 + 27x + 60\]\[=3(x^2+9x+20)\]\[=3(x+...)(x+...)\] the missing terms are the numbers that add to give nine and multiply to give 20